Determine the electric field and the potential at any point in space produced by a spherical


crown where the internal radius is R1 and the external one R2, with a total charge Q, for the


following cases: (a) non conducting and an uniform charge distributed throughout the volume;


and (b) metallic and on electrostatic equilibrium. Particularize the results for a solid sphere with radius R

1
Expert's answer
2022-04-04T09:07:27-0400

Answer

a) for non conducting sphere

Inside the sphere

"E=\frac{KQr}{R^3}"

At surface

"E=\frac{KQ}{R^2}"

Outside

"E=\frac{KQ}{r^2}"

Now potential

Inside to at surface remains same

"V=\frac{KQ}{R}"

Outside

"V=\frac{KQ}{r}"

b) now for matellic sphere

Inside E=0

Surface


"E=\frac{KQ}{R^2}"

At outside

"E=\frac{KQ}{r^2}"

Potential

Inside to surface remains same

"V=\frac{KQ}{R}"


Outside


"V=\frac{KQ}{r}"